# Visual Planning: Let’s Think Only with Images

### v1

@@ C.3 Out-of-Distribution Performance
Figure 9 illustrates generated images from VPFT and VPRL on OOD scenarios across Maze, FrozenLake, and MiniBehavior tasks. Notably, both models exhibit certain level of visual generalization to unseen configurations, such as larger grids with finer step granularity, despite not encountering them during training.
We subsequently quantitatively test generalization by evaluating the model on OOD environments with larger grid sizes. We find that SFT models performs poorly, while VPRL still demonstrates a certain level of visual planning capability as shown in Table 7. VPRL consistently outperforms VPFT in both Exact Match and Progress Rate, suggesting that it, to some degree, captures underlying planning strategies rather than merely memorizing training patterns.
C.4 Ablation: The Role of Stage 1
Exact Match (%)
Model 3$\times$3 4$\times$4 5$\times$5 6$\times$6
VPFT* 86.4 73.6 50.0 33.2
VPFT 92.0 82.8 68.8 58.0
@@ C.4 Ablation: The Role of Stage 1
Model FrozenLake (7$\times$7) Maze (7$\times$7) MiniBehavior (9$\times$9) EM (%) PR (%) EM (%) PR (%) EM (%) PR (%) VPFT 9.6 15.3 9.2 17.8 0.0 5.8 VPRL 20.4 31.2 10.0 21.6 0.4 14.7
C.5 Visual Planning Results
@@ Appendix D Prompting Templates
<image>
@@
Provide your final answer enclosed between <ANSWER> and </ANSWER>, for example: <ANSWER>right up
<image>

### v2

@@ C.4 Out-of-Distribution Performance
Figure 11 illustrates generated images from VPFT and VPRL on OOD scenarios across Maze, FrozenLake, and MiniBehavior tasks. Notably, both models exhibit a certain level of visual generalization to unseen configurations, such as larger grids with finer step granularity, despite not encountering them during training.
We subsequently quantitatively test generalization by evaluating the model on OOD environments with larger grid sizes. We find that SFT models perform poorly, while VPRL still demonstrates a certain level of visual planning capability, as shown in Table 9. VPRL consistently outperforms VPFT in both Exact Match and Progress Rate, suggesting that it, to some degree, captures underlying planning strategies rather than merely memorizing training patterns.
Finally, we analyze the robustness of VPRL by qualitatively testing its behavior under perturbed inputs. As shown in Figure 12, we mask portions of the input images with black or gray patches to simulate partial occlusion of the environment. Remarkably, the model continues to produce coherent planning traces within the masked environments, while preserving structural consistency with the visible input regions. This observation highlights the generalization capability of our visual planner, as it adapts to incomplete visual information without deviating from the underlying environment layout.
Figure 12: Qualitative analysis of VPRL under perturbed inputs (the first image of each trace). When parts of the input environment are masked (black/gray regions), VPRL maintains consistent planning traces aligned with the visible structure, demonstrating robustness to incomplete visual information without deviating from the underlying environment layout.
C.5 Ablation: The Role of Stage 1
@@
C.6 Visual Planning Results
@@ Appendix D Prompting Templates
FrozenLake (Coordinate & ASCII Representation)
Task: Frozen Lake Shortest Path Planning
- An elf marks the starting position.
- A gift represents the goal.
- Some cells contain ice holes that are impassable for the elf.
Describe the layout of the environment based on your analysis of the image, then provide your final answer enclosed between <ANSWER> and </ANSWER>, for example: <ANSWER>right up up</ANSWER>.
FrozenLake (GRPO)
Task: Frozen Lake Shortest Path Planning
- An elf marks the starting position.
- A gift represents the goal.
- Some cells contain ice holes that are impassable for the elf.
Present your reasoning enclosed within <think> and </think> tags. For example:
<think>Reasoning steps go here.</think>
Then, provide your final answer enclosed within <answer> and </answer> tags. For example:
<answer>right up up</answer>